# Unlocking new dimensions in photonic computing using optical skyrmions

### v4

@@
Unlocking new dimensions in photonic computing using optical skyrmions
@@
An Aloysius Wang Department of Engineering Science, University of Oxford, Parks Road, Oxford, OX1 3PJ, UK Corresponding authors: aw6609@princeton.edu, chao.he@eng.ox.ac.uk Yifei Ma Department of Engineering Science, University of Oxford, Parks Road, Oxford, OX1 3PJ, UK Yunqi Zhang Department of Engineering Science, University of Oxford, Parks Road, Oxford, OX1 3PJ, UK Zimo Zhao Department of Engineering Science, University of Oxford, Parks Road, Oxford, OX1 3PJ, UK Yuxi Cai Department of Engineering Science, University of Oxford, Parks Road, Oxford, OX1 3PJ, UK Xuke Qiu Department of Engineering Science, University of Oxford, Parks Road, Oxford, OX1 3PJ, UK Bowei Dong Institute of Microelectronics (IME), Agency for Science, Technology and Research (A*STAR), 2 Fusionopolis Way, Innovis #08-02, Singapore 138634, Republic of Singapore Chao He Department of Engineering Science, University of Oxford, Parks Road, Oxford, OX1 3PJ, UK Corresponding authors: aw6609@princeton.edu, chao.he@eng.ox.ac.uk
@@
The decline of Moore’s law coupled with the growing dominance of artificial intelligence has recently motivated research into photonic computing as a high-bandwidth, low-power strategy to accelerate digital electronics. However, modern-day photonic computing strategies are predominantly analog, making them susceptible to noise and intrinsically difficult to scale. The optical skyrmion offers a route to overcoming these limitations through digitization in the form of a discrete topological number that can be assigned to the analog optical field. Apart from an intrinsic robustness against perturbations, optical skyrmions represent a new dimension that has yet to be exploited for photonic computing, namely spatially varying polarization. Here, we propose a method of performing perturbation-resilient integer arithmetic with this new dimension of optical skyrmions through a family of passive optical components and present experimental evidence demonstrating its feasibility. To the best of our knowledge, this is the first time such discrete mathematical operations have been directly achieved using optical skyrmions without external energy input.
@@
Recent developments in structured light have enabled the generation of optical skyrmions [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19], which are complex spatially varying polarization fields that carry information through a topological number taking values in the integers. Optical skyrmions have three crucial properties that make them ideal candidates for high-density data transfer [15], namely the ability to interface with digital information given the discrete nature of the skyrmion number, a hypothesized robustness to perturbations, and the potential to store arbitrarily large integers within a single localized analog optical field.
@@
These same reasons also make the optical skyrmion a natural candidate for computation, with its integer-valued topological number offering a route to digital photonic computing in a way that transcends the usual bitwise framework of digital electronics. Moreover, compared to existing photonic computing strategies, which predominantly modulate amplitude, phase, and wavelength, complex spatially varying polarization fields represent an untapped dimension that can be independently manipulated and, therefore, has the potential to increase bandwidth significantly.
@@
Fig. 1: Concept. a, An optical skyrmion of arbitrary order $n$ passing through specially designed passive structured matter can have the effect of addition or subtraction by an arbitrary integer, $n\mapsto n\pm k$. Note that the same medium can be used to perform both addition and subtraction. Such structured matter can be realized in many different ways including SLM cascades, metasurfaces, inkjet printing of liquid crystal structures, gradient index systems, direct laser writing of birefringent structures in silica, and more. This figure also depicts the Stokes fields of a standard Néel-type skyrmion passing through such a medium. Throughout this paper, color is used to represent azimuthal angle on the Poincaré sphere, and saturation to represent height (similar to [15]). b, Stokes fields of skyrmions passing through different adders of first, second, and third order, and their respective skyrmion numbers. Here, we use the example of a linear retarder array described in the main text. For clarity, the only material property shown explicitly is the local axis of the retarder. Half waveplates that control addition and subtraction are also shown (see main text for details). c, Adders of first, second, and third order using highly disordered materials, demonstrating the robustness of our proposed adder to imperfections of the medium. Here, we consider perturbations to the linear retarder arrays proposed in the main text, which result in ellipticity of the axes as shown in the figure.
@@
The central reason for exploring the use of optical skyrmions in photonic computing is, however, their hypothesized resistance to noise and perturbations which arises from the integer-valued nature of the skyrmion number. The topological robustness of skyrmions has already been established in various domains, including magnetic skyrmions [20, 21, 22], and more recently, non-local quantum skyrmions [23, 24], where the idea of using skyrmions to digitize information is developed from a different perspective. In the setting of optical skyrmions carried through polarization fields, the study of topological robustness has also recently garnered attention [25], with a degree of protection already verified for propagation through a number of different media [26, 10, 14]. Robustness to noise is particularly relevant to photonic computing, which has thus far remained analog and, therefore, susceptible to noise. Take, for instance, the following three common photonic accelerators. The accumulation of random phase errors in Mach-Zehnder interferometer meshes [27, 28], thermal crosstalk in microring resonator weight banks [29, 30] and low optical contrast in phase-change material-based photonic crossbar arrays [31] all lead to degraded signal-to-noise ratio in large-size implementations, circumventing the scalability of these architectures. Even with progress in device innovation [32, 33], system calibration, and control algorithm optimization to minimize errors[34, 35], almost all demonstrated photonic accelerators thus far have had a size of less than 4 inputs by 4 outputs. Computing using optical skyrmions is, therefore, a promising solution offering greater scalability through improved stability against noise.
@@
Motivated by this possibility, we describe here a class of structured matter (see Methods 1) which, when restricted to particular types of skyrmions, behave as adders/subtractors (Fig. 1a) and provide experimental evidence supporting this fact. Note that the same medium can perform both addition and subtraction, but we adopt the term “adder” throughout this paper for brevity. We would also like to emphasize that skyrmions are not topologically protected through the types of structured matter described in this paper, which have been specially designed to manipulate the skyrmion number. However, the topological nature of the skyrmion gives rise to resilience against perturbations in a different form, namely that material parameters have the flexibility to fluctuate without affecting the medium’s ability to perform arithmetic (see Methods 1). As mentioned above, this resilience of function is of key importance in photonic computing. Such structured matter can be realized using continuously varying retarders with specific structures on their boundary and offer a reliable way of manipulating the skyrmion number without external energy input. Moreover, we demonstrate that by adopting a more generalized skyrmion number [36], restrictions at the boundary can be further loosened at the cost of increased computational complexity. Our work opens the doors to this entirely new framework for performing photonic computing, where optical skyrmions are used as the fundamental unit in computation.
@@ 1 Main
Beginning with conventional skyrmion adders, we provide a precise description of our proposed optical component, followed by experimental results that demonstrate its feasibility and robustness to disorder. As detailed in Methods 1, given a general homotopy of skyrmions, the resulting difference in skyrmion number is equal to the skyrmion number of the homotopy when restricted onto the walls of the homotoping cylinder. This transformation law gives a systematic way of engineering structured matter to perform addition.
For example, consider a spatially varying retarder [37], which when restricted to its boundary, is linear and has a retardance of $\pi$. Let $k$ be the number of half-revolutions made by the axis of the retarder traversing counterclockwise along its boundary (Fig. 1b). Then, with no further restrictions to material properties apart from continuity, one may show
$\deg\mathcal{S}^{\prime}=\left\{\begin{array}[]{r l}\deg\mathcal{S}+k&\text{if$\mathcal{S}$ is right circularly polarized (RCP) on its boundary}\\ \deg\mathcal{S}-k&\text{if $\mathcal{S}$ is left circularly polarized (LCP) onits boundary}\end{array}\right.$ (1)
@@
Fig. 2: Adder modules and experimental results. a, Order $k$ adder modules can be constructed using a linear retarder array placed before a half waveplate while order $k$ subtractor modules can be constructed using a linear retarder array placed after a half waveplate. The addition of the half waveplates resolves incompatibilities in polarization states on the boundary as explained in the main text, allowing for the different modules to be cascaded indefinitely. Notice also that order $k$ adders and subtractors can be realized using the same hardware, with one direction performing addition and the other performing subtraction. Lastly, while we use gradient index systems and a 3 SLM cascade to implement the adder in our work, it is worth emphasizing that an adder can actually be implemented in many different ways, as long as its properties at the boundary are properly constrained. b, A subset of the measured Stokes fields of optical skyrmions passing through adders of order 2 realized using gradient index systems. The operations $1\pm 2\pm 2$ and $3\pm 2\pm 2$ are shown with detailed implementation presented in Supplementary Note 1 and the full dataset presented in Supplementary Fig. 2. c, A subset of the measured Stokes fields of optical skyrmions passing through adders of order 1 and 3 realized using a 3 SLM cascade, and where disorder is simulated by a random pixel-wise noise to the voltage levels of the SLMs. Three levels of disorder are shown as indicated by the color of the star, increasing from left to right. Details of the implementation can be found in Supplementary Note 1, and a complete dataset is presented in Supplementary Fig. 3.
@@
Our mathematical results also imply that the function of any medium designed in this way depends only on the structure of its boundary and is independent of material properties everywhere else provided they are continuous (excluding certain extreme cases; see Discussion and Methods 1). This is a reflection of the topological structure of the optical skyrmion exhibited in matter. From a practical perspective, this suggests a strong robustness of any such adder to physical imperfections of the medium that implements it, which greatly eases fabrication. We demonstrate the robustness arising from this “topological duality” between field and matter in Fig. 1c, which shows the invariance of the effect of our proposed adder on the skyrmion number of optical fields to perturbations in the medium, and give a more detailed explanation of robustness and duality in Methods 1. Note also that while other works [9, 19] have investigated similar structured linear retarder arrays, our work demonstrates, for the first time, the topological properties of such structured matter, and their ability to modify the topological number of non-uniform input fields.
@@
After passing through the adder proposed above, the boundary conditions of the skyrmion will flip, that is, RCP becomes LCP and vice versa. Therefore, to cascade multiple adders together, one must realign the outer boundary after each operation, which is most easily done using a homogenous waveplate. With our proposed design, this amounts to adding a half waveplate after the medium for addition and a half waveplate before for subtraction, provided we design our components for skyrmions that are RCP on their boundary (Fig. 2a and Supplementary Fig. 6). Therefore, combining an adder with a half waveplate effectively produces a cascadable module that performs addition in one direction and subtraction in the other.
@@
Next, to demonstrate the feasibility of our proposed medium, we provide experimentally measured Stokes fields of skyrmions passing through second-order adders realized via gradient index systems. Fig. 2b shows a subset of our results, with the full dataset presented in Supplementary Fig. 2. We generated skyrmions of orders ranging from $-3$ to 3 using a cascade of 2 spatial light modulators (SLMs) and passed the field through the medium with appropriate waveplates to achieve the operations $+2+2$, $+2-2$, $-2+2$ and $-2-2$ (see Supplementary Note 1 for details of the techniques used in beam generation, measurement and analysis, and Supplementary Fig. 1 for the experimental assembly). Note the medium we used was discarded by quality assurance due to its asymmetric axis distribution, and it exhibits significant perturbations in material properties compared to the typical sample. Despite this, the numerically computed skyrmion numbers show that our proposed adder efficiently and reliably performs the desired operations, even with imperfections in the medium. Further technical details of the experiments are presented in Supplementary Note 1.
@@
Fig. 3: Generalized skyrmion adders. a, Concept of a generalized skyrmion photo-adder, which is a passive component that converts a skyrmion of degree $n$ into a generalized skyrmion of degree $(n+k_{1},n+k_{2},n+k_{3},\ldots)$. Note that the function of the adder is robust to perturbations in both the input field and material parameters, with this robustness extending even to situations where perturbations occur at the boundary. b, Given a polarization field, a single generalized skyrmion number can be defined for each connected component of the Poincaré sphere (PS) carved out by the image of the boundary curve. A field with (left) one component and (right) three components, along with the corresponding images of their boundary curves on the Poincaré sphere are shown. A stereographically projected version of the boundary curve is also shown. Note that for a given boundary condition, any continuous extension of the boundary to the entire domain will have the same number of generalized skyrmion numbers. c, A generalized skyrmion adder works by manipulating the boundary to create new connected components. For each newly created component, the original skyrmion number is increased once for each time the boundary curve encircles the component, accounting for orientation. The figure depicts an example of a $(n)\mapsto(n+1,n-1,n)$ adder, with input field $n=2$ and where the stokes fields and stereographically projected boundary curves are shown. Finally, the skyrmion number and generalized skyrmion numbers of the two fields are provided.
@@
To further emphasize the perturbation resilience of our proposed adder and the versatility of its implementation, we present experimental results where a cascade of 3 SLMs is used to realize an adder and disorder is simulated by introducing random pixel-wise noise to the voltage levels of the SLMs. Note that the tunability of the SLMs is used to introduce disorder of varying strengths, but it is not central to how the adder operates (that is, the SLM cascade merely mimics a passive device). The noise is added in such a way that it is maximum at the center and gradually decreases to zero at the boundary. Fig. 2c shows a subset of our results with the full dataset, including Mueller matrices of the disordered media and polarization ellipses, presented in Supplementary Fig. 3. Observe from the figure that the disorder we have added is significantly larger than what would typically occur in practice. Additionally, systematic errors due to phase unwrapping lead to lines observed in the output Stokes field, and this can also be considered a form of perturbation. Nonetheless, the numerically computed skyrmion number remains stable, demonstrating the strong topological robustness of our adder.
@@
Lastly, we describe a new method to simultaneously enhance the topological robustness of our proposed adder against perturbations in the state of polarization (SoP) of the input light at the boundary, as well as against perturbations in the material parameters at the boundary, by adopting the generalized skyrmion number introduced in [36]. The generalized skyrmion number is a method of assigning non-compactifiable fields [26] a tuple of integers $(n_{1}\ldots,n_{k})$ which are topologically protected under a general notion of homotopy described in [36]. Note that this represents an increase in the dimension of the information carried by the field, thereby significantly improving its information density.
@@
On a more technical level, given any smooth polarization field $\mathcal{S}\colon\Omega\longrightarrow S^{2}$, we can define one generalized skyrmion number for each connected component of the Poincaré sphere carved out by the image of the boundary curve $\mathcal{S}\lvert_{\partial\Omega}$ (Fig. 3b). The generalized skyrmion number associated with a connected component can be computed by the integral equation
[…]
@@
Specializing to our proposed adder (whose axis distribution is as depicted in Fig. 1b), if the outer retardance is a constant that lies between 00 and $\pi$, three cases arise depending on the SoP at the boundary of the incident skyrmion. For SoPs close to RCP, one has the transition $(n)\mapsto(n+k,n)$, for SoPs close to LCP, one has the transition $(n)\mapsto(n-k,n)$, and for all other states, one has the transition $(n)\mapsto(n+k,n-k,n)$. Thus, one can essentially select the adder’s function by using different input boundary SoPs.
Notice also that by adopting the generalized skyrmion number, not only can we have situations where multiple additions and subtractions occur simultaneously, but there is also a general tolerance to the SoP of the input beam. Indeed, the adder splits the Poincaré sphere into various regions, and performs a specific type of operation on each region. More generally, as long as the number of new connected components that form during propagation and the orientation in which the boundary curve encircles each component remains the same, the same operations will be observed at the level of the generalized skyrmion number. This latter property also makes the generalized skyrmion adder stable against fluctuations in the parameters of the matter, without the strict constraints at the boundary present in the skyrmion adders introduced earlier.
Fig. 4: Experimental results (generalized skyrmion adders). a, (Left) The different regions of the Poincaré sphere on which the function of our adder is stable. The top region corresponds to addition, the bottom to subtraction and the middle to both addition and subtraction simultaneously. (Right) A subset of the measured output stokes fields and computed generalized skyrmion numbers at different levels of disorder (increasing from left to right) and different incident SoPs picked to demonstrate addition, subtraction and addition and subtraction simultaneously. A stereographic projection of the boundary curve is also shown, with the change in skyrmion number of each region labeled. Note the color of these labels indicate the generalized skyrmion number corresponding to that region. A complete dataset including details of the incident SoPs and measured Mueller matrices is presented in Supplementary Fig. 4 . Technical details such as the formation of small loops due to disorder are also discussed. Lastly, details of the implementation and the experimental assembly can be found in Supplementary Note 1 and Supplementary Fig. 1, respectively. b, The skyrmion number and generalized skyrmion numbers at each level of disorder for different incident SoPs. Note that the generalized skyrmion number is topologically stable even though the usual skyrmion number is not.
@@
Moreover, in the full dataset presented in Supplementary Fig. 4, we demonstrate that the function of the adder is stable within a range of SoPs. This stability of the generalized skyrmion adder to different inputs is a important feature which is not enjoyed by the skyrmion adders introduced earlier. Lastly, a comparison between the skyrmion number and generalized skyrmion numbers at different levels of disorders is also shown in Fig. 4b. From the plots, it is clear that the generalized skyrmion number is topologically stable even though the usual skyrmion number is not. The technique of computing the generalized skyrmion number and estimating the boundary curve is adapted from [36].

### v6

# Perturbation-resilient integer arithmetic using optical skyrmions

@@
Perturbation-resilient integer arithmetic using optical skyrmions
@@
An Aloysius Wang Department of Engineering Science, University of Oxford, Parks Road, Oxford, OX1 3PJ, UK Corresponding authors: aloysius.wang@gmail.com, chao.he@eng.ox.ac.uk Yifei Ma Department of Engineering Science, University of Oxford, Parks Road, Oxford, OX1 3PJ, UK Yunqi Zhang Department of Engineering Science, University of Oxford, Parks Road, Oxford, OX1 3PJ, UK Zimo Zhao Department of Engineering Science, University of Oxford, Parks Road, Oxford, OX1 3PJ, UK Yuxi Cai Department of Engineering Science, University of Oxford, Parks Road, Oxford, OX1 3PJ, UK Xuke Qiu Department of Engineering Science, University of Oxford, Parks Road, Oxford, OX1 3PJ, UK Bowei Dong Institute of Microelectronics (IME), Agency for Science, Technology and Research (A*STAR), 2 Fusionopolis Way, Innovis #08-02, Singapore 138634, Republic of Singapore Chao He Department of Engineering Science, University of Oxford, Parks Road, Oxford, OX1 3PJ, UK Corresponding authors: aloysius.wang@gmail.com, chao.he@eng.ox.ac.uk
@@
The decline of Moore’s law coupled with the rise of artificial intelligence has recently motivated research into photonic computing as a high-bandwidth, low-power strategy to accelerate digital electronics. However, many modern-day photonic computing strategies are analog, making them susceptible to noise and intrinsically difficult to scale. Optical skyrmions offer a route to overcoming these limitations through digitization in the form of a discrete topological number that can be assigned to the analog optical field. Apart from an intrinsic robustness against perturbations, optical skyrmions represent a new medium that has yet to be fully exploited for photonic computing, namely spatially varying polarization. Here, we propose and experimentally demonstrate a method for performing perturbation-resilient integer arithmetic with optical skyrmions and passive optical components. To the best of our knowledge, this is the first time such discrete mathematical operations have been directly achieved using optical skyrmions without external energy input.
@@
Recent developments in structured light have enabled the generation of optical skyrmions [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], which include complex spatially varying polarization fields that carry information through a topological number taking values in the integers. Optical skyrmions have three crucial properties that make them ideal candidates for high-density data transfer [13], namely the ability to interface with digital information given the discrete nature of the skyrmion number, a robustness to perturbations, and the potential to store arbitrarily large integers within a single localized analog optical field.
@@
These same reasons also make optical skyrmions (here referring specifically to Stokes skyrmions) a natural candidate for computation, with its integer-valued topological number offering a route to digital photonic computing in a way that transcends the usual bitwise framework of digital electronics. Moreover, compared to existing photonic computing strategies, which predominantly modulate amplitude, phase, and wavelength, complex spatially varying polarization fields represent an untapped dimension that can be independently manipulated and, therefore, has the potential to increase bandwidth significantly. This is especially relevant given the growing recognition of untapped spatial degrees of freedom as a means to advance photonic technologies, with free-space optical skyrmions being one such approach [19, 20].
@@
Fig. 1: Concept. a, An optical skyrmion of arbitrary order $n$ passing through specially designed passive structured matter can have the effect of addition or subtraction by an arbitrary integer, $n\mapsto n\pm k$ (only $n\mapsto n+k$ is shown here). Note that the same medium can be used to perform both addition and subtraction. Such structured matter can be realized in many different ways including spatial light modulator (SLM) cascades, metasurfaces, inkjet printing of liquid crystal structures, gradient index systems, direct laser writing of birefringent structures in silica, and more. This figure also depicts the Stokes fields of a standard Néel-type skyrmion passing through such a medium. Throughout this paper, color is used to represent azimuthal angle on the Poincaré sphere (PS), and saturation to represent height (similar to [13]). b, Stokes fields of skyrmions passing through different adders of first, second, and third order, and their respective skyrmion numbers. The linear retarder array described in the main text is used as an example, with spatially varying material properties illustrated using cylinders, where the local axis orientation determines the shape and color, and local retardance determines the height. Half-wave plates (HWPs) that control addition and subtraction are also shown (see main text for details). c, Adders of first, second, and third order using highly disordered materials, demonstrating the robustness of our proposed adder to imperfections of the medium. Here, we consider perturbations to the linear retarder arrays that respect the conditions proposed in the main text, resulting in ellipticity of the axes and changes in retardance. Note that the chosen disorder here is merely an example that abstractly represents an arbitrary level of distortion.
@@
The central reason for exploring the use of optical skyrmions in photonic computing is, however, their resistance to noise and perturbations which arises from the integer-valued nature of the skyrmion number. The topological robustness of skyrmions has already been established in various domains, including magnetic skyrmions [21, 22, 23], and more recently, non-local quantum skyrmions [24, 25], where the idea of using skyrmions to digitize information is developed from a different perspective. In the setting of optical skyrmions carried through polarization fields, the study of topological robustness has also recently garnered attention [26], with a degree of protection already verified for propagation through a number of different media [27, 8, 12]. This robustness is particularly relevant to photonic computing, which has thus far remained analog and, therefore, susceptible to noise. Take, for instance, the following three common photonic accelerators. The accumulation of random phase errors in Mach-Zehnder interferometer meshes [28, 29], thermal crosstalk in microring resonator weight banks [30, 31] and low optical contrast in phase-change material-based photonic crossbar arrays [32] all lead to degraded signal-to-noise ratio in large-size implementations, circumventing the scalability of these architectures. Even with progress in device innovation [33, 34], system calibration, and control algorithm optimization to minimize errors[35, 36], almost all photonic accelerators demonstrated thus far have remained small in size (typically less than 4 inputs by 4 outputs). While recent breakthroughs have enabled larger-scale architectures [37, 38], it has been noted that noise remains a key challenge and a limiting factor for high-throughput data manipulation in photonic chips due to the use of analog photonic signals in computing. Lastly, while spatial degrees of freedom have been utilized for computation-related applications through orbital angular momentum before [39, 40, 41], these approaches do not exhibit the same robustness to noise as they are generally linear-algebraic rather than topological in nature. Computing using optical skyrmions is, therefore, a promising solution offering greater scalability through improved stability against noise.
@@
Motivated by this possibility, we describe here a class of structured matter (see Methods 1) which, when restricted to particular types of input skyrmions, behaves as adders/subtractors (Fig. 1a) and provide experimental evidence supporting this fact. Note that the same medium can perform both addition and subtraction, but we adopt the term “adder” throughout this paper for brevity. We would also like to emphasize that skyrmions are not topologically protected through the types of structured matter described in this paper, which have been specially designed to manipulate the skyrmion number. However, the topological nature of the skyrmion gives rise to resilience against perturbations in a different form, namely that material parameters have the flexibility to fluctuate without affecting the medium’s ability to perform arithmetic (see Methods 1). As mentioned above, this resilience of function is of key importance in photonic computing. Such structured matter can be realized using continuously varying retarders with specific structures on their boundary and offer a reliable way of manipulating the skyrmion number without external energy input. Moreover, we demonstrate that by adopting a generalized skyrmion number [42], it is possible to simultaneously increase the dimensionality of information carried while relaxing boundary restrictions. Our work opens the doors to this entirely new framework for performing photonic computing, where optical skyrmions are used as the fundamental unit in computation.
@@ 1 Main
Beginning with conventional skyrmion adders, we provide a mathematical description of our proposed optical component, followed by experimental results that demonstrate its feasibility and robustness to disorder. As detailed in Methods 1, given a general homotopy of skyrmions, the resulting difference in skyrmion number is equal to the skyrmion number of the homotopy when restricted onto the walls of the homotoping cylinder. This transformation law gives a systematic way of engineering structured matter to perform addition.
For example, consider a spatially varying retarder, which when restricted to its boundary, is linear and has a retardance of $\pi$. Let $k$ be the number of half-revolutions made by the axis of the retarder traversing counterclockwise along its boundary (Fig. 1b). Then, with no further restrictions to material properties apart from continuity, one may show
[…]
@@
Fig. 2: Adder modules and experimental results. a, Order $k$ adder modules can be constructed using a linear retarder array placed before a half-wave plate while order $k$ subtractor modules can be constructed using a linear retarder array placed after a half-wave plate. The addition of the half-wave plates resolves incompatibilities in polarization states on the boundary as explained in the main text, allowing for the different modules to be cascaded indefinitely. Notice also that order $k$ adders and subtractors can be realized using the same hardware, with one direction performing addition and the other performing subtraction. Lastly, while gradient index systems and a 3-SLM cascade are used to implement the adder in our work, it is worth emphasizing that an adder can be implemented in many different ways, provided its properties at the boundary are properly constrained. b, A subset of the measured Stokes fields of optical skyrmions passing through adders of order 2 realized using gradient index systems. The operations $1\pm 2\pm 2$ and $3\pm 2\pm 2$ are shown with detailed implementation presented in Supplementary Note 1 and the full dataset presented in Supplementary Fig. 2. c, A subset of the measured Stokes fields of optical skyrmions passing through adders of order 1 and 3 realized using a 3 SLM cascade, and where disorder is introduced by a random pixel-wise noise to the voltage levels of the SLMs. Since only a single adder is used in this experiment, no half-wave plate is included. Three levels of disorder are shown as indicated by the color of the star, increasing from left to right. Details of the implementation can be found in Supplementary Note 1, and a complete dataset is presented in Supplementary Fig. 3.
@@
Our mathematical results also imply that the function of any medium designed in this way depends only on the structure of its boundary and is independent of material properties everywhere else provided they are continuous (excluding certain extreme cases; see Discussion and Methods 1). This is a reflection of the topological structure of the optical skyrmion exhibited in matter. From a practical perspective, this suggests a strong robustness of any such adder to physical imperfections of the medium that implements it, which greatly eases fabrication. We demonstrate the robustness arising from this “topological duality” between field and matter in Fig. 1c, which shows the invariance of the effect of our proposed adder on the skyrmion number of optical fields to perturbations in the medium, and give a more detailed explanation of robustness and duality in Methods 1.
@@
After passing through the adder proposed above, the boundary conditions of the skyrmion will flip, that is, RCP becomes LCP and vice versa. Therefore, to cascade multiple adders together, one must realign the outer boundary after each operation. With our proposed design, this amounts to adding a half-wave plate after the medium for addition and a half-wave plate before for subtraction, provided we design our components for skyrmions that are RCP on their boundary (Fig. 2a and Supplementary Fig. 6). Therefore, combining an adder with a half-wave plate effectively produces a cascadable module that performs addition in one direction and subtraction in the other.
@@
Next, to demonstrate the feasibility of our proposed medium, we provide experimentally measured Stokes fields of skyrmions passing through second-order adders realized via gradient index systems. Fig. 2b shows a subset of our results, with the full dataset presented in Supplementary Fig. 2. We generated skyrmions of orders ranging from 3-3 to 3 using a cascade of 2 SLMs and passed the field through the medium with appropriate waveplates to achieve the operations +2+2+2+2, +22+2-2, 2+2-2+2 and 22-2-2 (see Supplementary Note 1 for details of the techniques used in beam generation, measurement and analysis, and Supplementary Fig. 1 for the experimental assembly). Note the medium we used was discarded by quality assurance due to its asymmetric axis distribution, and it exhibits significant perturbations in material properties compared to the typical sample. Despite this, the numerically computed skyrmion numbers show that our proposed adder efficiently and reliably performs the desired operations, even with imperfections in the medium. Further technical details of the experiments are presented in Supplementary Note 1.
@@
Fig. 3: Generalized skyrmion adders. a, Concept of a generalized skyrmion photo-adder, which is a passive component that converts a skyrmion of degree $n$ into a generalized skyrmion of degree $(n+k_{1},n+k_{2},n+k_{3},\ldots)$. Note that the function of the adder is robust to perturbations in both the input field and material parameters, with this robustness extending even to situations where perturbations occur at the boundary. b, Given a polarization field, a single generalized skyrmion number can be defined for each connected component of the Poincaré sphere carved out by the image of the boundary curve. A field with (left) one component and (right) three components, along with the corresponding images of their boundary curves on the Poincaré sphere are shown. A stereographically projected version of the boundary curve is also shown. Note that for a given boundary condition, any continuous extension of the boundary to the entire domain will have the same number of generalized skyrmion numbers. c, A generalized skyrmion adder works by manipulating the boundary to create new connected components. For each newly created component, the original skyrmion number is increased once for each time the boundary curve encircles the component, accounting for orientation. The figure depicts an example of a $(n)\mapsto(n+1,n-1,n)$ adder, with input field $n=2$ and where the Stokes fields and stereographically projected boundary curves are shown. Finally, the skyrmion number and generalized skyrmion numbers of the two fields are provided.
@@
To further emphasize the perturbation resilience of our proposed adder and the versatility of its implementation, we present experimental results where a cascade of 3 SLMs is used to realize an adder and disorder is simulated by introducing random pixel-wise noise to the voltage levels of the SLMs. Since the 3 SLM cascade is designed to achieve arbitrary retardance and axis orientation [8], adding noise to the voltage levels of the SLMs effectively simulates a disordered array of arbitrary elliptical retarders, which has the added benefit of reflecting a wide variety of real-world perturbations. Note that the tunability of the SLMs is used to introduce disturbances of varying strengths, but it is not central to how the adder operates (that is, the SLM cascade merely mimics a passive device). The noise is added in such a way that it is maximum at the center and gradually decreases to zero at the boundary, consistent with the derivation in Methods 1. Fig. 2c shows a subset of our results with the full dataset, including Mueller matrices of the disordered media and polarization ellipses, presented in Supplementary Fig. 3. Observe from the figure that the disorder we have added is significantly larger than what would typically occur in practice. Additionally, systematic errors due to phase unwrapping lead to lines observed in the output Stokes field, and this can also be considered a form of perturbation. Nonetheless, the numerically computed skyrmion number remains stable, demonstrating the strong topological robustness of our adder. There remains significant scope for further exploration into the limits of topological protection of optical skyrmions in the presence of random noise, including the effects of spatially correlated noise and the limits at which topological protection breaks down (see details in Supplementary Note 3), which we plan to address in future work.
@@
Lastly, by adopting the generalized skyrmion number introduced in [42], it is possible to simultaneously enhance the topological robustness of our proposed adder against perturbations in the state of polarization (SoP) of the input light at the boundary, as well as against perturbations in the material parameters at the boundary. Moreover, this approach allows a single field to carry multiple topological charges, representing an increase in the dimensionality of the information carried by the field and significantly improving its information density.
@@
On a more technical level, the generalized skyrmion number is a method of assigning non-compactifiable fields [27] a tuple of integers $(n_{1}\ldots,n_{k})$ derived from the De Rham cohomology of compactly supported forms which are topologically protected under a general notion of homotopy [42]. Given any smooth polarization field $\mathcal{S}\colon\Omega\longrightarrow S^{2}$, we can define one generalized skyrmion number for each connected component of the Poincaré sphere carved out by the image of the boundary curve $\mathcal{S}\lvert_{\partial\Omega}$ (Fig. 3b). The generalized skyrmion number associated with a connected component can be computed by the integral equation
$\text{generalized skyrmion number}=\frac{1}{c}\int_{\Omega}f(S)S\cdot\left(\frac{\partial S}{\partial x}\times\frac{\partial S}{\partial y}\right)dxdy,$
@@
Specializing to our proposed adder (whose axis distribution is as depicted in Fig. 1b), if the outer retardance is a constant that lies between 0 and $\pi$, three cases arise depending on the SoP at the boundary of the incident skyrmion. For SoPs close to RCP, one has the transition $(n)\mapsto(n+k,n)$, for SoPs close to LCP, one has the transition $(n)\mapsto(n-k,n)$, and for all other states, one has the transition $(n)\mapsto(n+k,n-k,n)$. Thus, one can essentially select the adder’s function by using different input boundary SoPs.
Notice also that by adopting the generalized skyrmion number, not only can we have situations where multiple additions and subtractions occur simultaneously, but there is also a general tolerance to the boundary SoP of the input beam. Indeed, the adder splits the Poincaré sphere into various regions, and performs a specific type of operation on each region. More generally, as long as the number of new connected components that form during propagation and the orientation in which the boundary curve encircles each component remains the same, the same operations will be observed at the level of the generalized skyrmion number. This latter property makes the generalized skyrmion adder stable against fluctuations in the parameters of the matter, without the strict constraints at the boundary present in the skyrmion adders introduced earlier. Lastly, note that while it is possible to manipulate fields with non-integer skyrmion numbers for computation, the skyrmion number in this case is not a topologically protected quantity, and therefore does not support the same level of robustness as the generalized skyrmion number (see details in Supplementary Note 2).
Fig. 4: Experimental results (generalized skyrmion adders). a, (Left) The different regions of the Poincaré sphere on which the function of our adder is stable. The top region corresponds to addition, the bottom to subtraction and the middle to both addition and subtraction simultaneously. (Right) A subset of the measured output Stokes fields and computed generalized skyrmion numbers is shown for different levels of disorder (increasing from left to right) and different incident SoPs selected to demonstrate addition, subtraction, and simultaneous addition and subtraction. A stereographic projection of the boundary curve is also shown, with the change in skyrmion number of each region labeled. Note the color of these labels indicate the generalized skyrmion number corresponding to that region. A complete dataset including details of the incident SoPs and measured Mueller matrices is presented in Supplementary Fig. 4 . Technical details such as the formation of small loops due to disorder are also discussed. Lastly, details of the implementation and the experimental assembly can be found in Supplementary Note 1 and Supplementary Fig. 1, respectively. b, The skyrmion number and generalized skyrmion numbers at each level of disorder for different incident SoPs. Note that the generalized skyrmion number is topologically stable even though the usual skyrmion number is not.
@@
Moreover, in the full dataset presented in Supplementary Fig. 4, we demonstrate that the function of the adder is stable within a range of SoPs. This stability of the generalized skyrmion adder to different inputs is an important feature which is not enjoyed by the skyrmion adders introduced earlier. Lastly, a comparison between the skyrmion number and generalized skyrmion numbers at different levels of disorders is also shown in Fig. 4b. From the plots, it is clear that the generalized skyrmion number is topologically stable even though the usual skyrmion number is not. The technique of computing the generalized skyrmion number and estimating the boundary curve is adapted from [42].